# Context-measure: Contextualizing Metric for Camouflage

### v2

@@ 3.1 Definition of Probabilistic Pixel Correlation
Let $X_{H\times W}$ and $Y_{H\times W}$ denote the predicted and GT masks, where $H$ and $W$ are the image height and width, respectively. For a pixel $p$, we define two foreground pixel sets: $X_{f}:=\{\text{pixel}\ p|X(p)>0\}$ and $Y_{f}:=\{\text{pixel}\ p|Y(p)>0\}$, where $X(p)\in[0,1]$ and $Y(p)\in\{0,1\}$ denote the values of pixel $p$. To quantify the correlation between any two pixels, we model their dependency using a probabilistic formulation. Specifically, we estimate the covariance matrix $\mathbf{\Sigma}_{Y}$ from $Y_{f}$, capturing the shape of the GT object. We normalize the covariance as $\hat{\mathbf{\Sigma}}_{Y}=\frac{\alpha^{2}}{\operatorname{Tr}(\mathbf{\Sigma}_{Y})}\mathbf{\Sigma}_{Y}$, where $\alpha$ is a scalar hyperparameter and $\operatorname{Tr}(\cdot)$ denotes the trace operator. This ensures that pixel-wise correlations remain comparable across images regardless of their resolution size. We define the correlation between two pixels conditioned on a specific foreground $Y_{f}$, which is written as:
@@ 4 Experiments & Analyses
Experimental Setup. All competing metrics were evaluated using their default configurations. In the following experiments, we analyses two variants of our measure. The generic variant, $C_{\beta}$ as defined in Eqn. 2, was implemented with $\alpha=6$ and $\beta=1.0$. The camouflaged variants, $C_{\beta}^{\omega}$ as in Eqn. 10, applied with a contextual band width $k=20$, patch size $N=7$ with an overlap of 3, and hyper-parameters $\alpha=6$, $\beta=1.2$, $\gamma=8$, and $\lambda=20$.
@@ Appendix A Revisiting Saliency-era Metrics
$\bullet$ F-measure ($F_{\beta}$). The F-measure [2, 3] is a standard evaluation metric that balances Precision and Recall when comparing the predicted mask with the ground-truth mask. Its computation depends on four commonly used pixel-level quantities: true positives ($TP$) — foreground pixels correctly predicted as foreground; true negatives ($TN$) — background pixels correctly predicted as background; false positives ($FP$) — background pixels incorrectly predicted as foreground; and false negatives ($FN$) — foreground pixels missed by the prediction. These quantities are computed as:
@@
where $\odot$ is element-wise multiplication, matrix $A$ models the spatial dependency between pixels, and vector $B$ encodes the varying perceptual importance across locations. Based on $E^{\omega}$, the weighted true positives, true negatives, false positives, and false negatives are computed as:

### v3

@@ 3.1 Definition of Probabilistic Pixel Correlation
Let $X_{H\times W}$ and $Y_{H\times W}$ denote the predicted and GT masks, where $H$ and $W$ are the image height and width, respectively. For a pixel $p$, we define two foreground pixel sets: $X_{f}:=\{\text{pixel}\ p|X(p)>0\}$ and $Y_{f}:=\{\text{pixel}\ p|Y(p)>0\}$, where $X(p)\in[0,1]$ and $Y(p)\in\{0,1\}$ denote the values of pixel pp. To quantify the correlation between any two pixels, we model their dependency using a probabilistic formulation. Specifically, we estimate the covariance matrix $\mathbf{\Sigma}_{Y}$ from $Y_{f}$, capturing the shape of the GT object. We normalize the covariance as $\hat{\mathbf{\Sigma}}_{Y}=\frac{\alpha^{2}}{\operatorname{Tr}(\mathbf{\Sigma}_{Y})}\mathbf{\Sigma}_{Y}$, where $\alpha$ is a scalar hyperparameter and $\operatorname{Tr}(\cdot)$ denotes the trace operator. This ensures that pixel-wise correlations remain comparable across images regardless of their resolution size. We define the correlation between two pixels conditioned on a specific foreground $Y_{f}$, which is written as:
@@ 4 Experiments & Analyses
Experimental Setup. All competing metrics were evaluated using their default configurations. In the following experiments, we analyses two variants of our measure. The generic variant, $C_{\beta}$ as defined in Eqn. 2, was implemented with $\alpha=6$ and $\beta=1.0$. The camouflaged variants, $C_{\beta}^{\omega}$ as in Eqn. 10, applied with a contextual band width $k=20$, patch size $N=7$ with an overlap of 3, and hyper-parameters $\alpha=6$, $\beta^{2}=1.2$, $\gamma=8$, and $\lambda=20$.
@@ Appendix A Revisiting Saliency-era Metrics
$\bullet$ F-measure ($F_{\beta}$). The F-measure [2, 3] is a standard evaluation metric that balances Precision and Recall when comparing the predicted mask with the ground-truth mask. Its computation depends on four commonly used pixel-level quantities: true positives (TPTP) — foreground pixels correctly predicted as foreground; true negatives (TNTN) — background pixels correctly predicted as background; false positives (FPFP) — background pixels incorrectly predicted as foreground; and false negatives (FNFN) — foreground pixels missed by the prediction. These quantities are computed as:
@@
where $\odot$ is element-wise multiplication, matrix AA models the spatial dependency between pixels, and vector $B$ encodes the varying perceptual importance across locations. Based on $E^{\omega}$, the weighted true positives, true negatives, false positives, and false negatives are computed as: